An electron in a hydrogen atom decays from a certain excited state down to the ground state. If the photon emitted has a wavelength of 97 nm, what was the n-value for the excited state?
E = hν = hc/λ
ΔE = RH(1/ni2 - 1/nf2)
RH = 2.18 x 10-18 J
Planck's constant = 6.63 x 10-34 J s
c = 3.00 x 108 m/s
1 nm = 10-9 m
1
Expert's answer
2020-11-09T14:10:28-0500

Note that energy will be negative since its losing an atom

ΔE=(hc)/(v×RH)= [ 1/ni2 - 1/nf2 ]

-(6 63×10-34 × 3.0×108 )/(97×10-9 ×2.18×10-18) = (1/12 - 1/nf2)

-0.94=1-(1/nf2)

-1.94= -1/nf2

nf2 = 1/1.94 = 0.515

nf = square root 0.515

nf= 0.72